OpenStudy (anonymous):

A perfectly spherical piece of metal is found at the bottom of a wishing well. The mass of the object is 1.3 kg and the radius is 0.25 m. What is its density? Answer in units of kg/m3. 009 (part 2 of 2) 10.0 points What would be its weight if it had the same volume and were made of pure gold? The density of pure gold is 19300 kg/m3. Answer in units of N.

OpenStudy (anonymous):

use density formula Density=Mass/volume the mass of it is 1.3k and the volume of it is .25 so 1.3/0.25= 5.2kg/m now cube that answer for final. For gold you have to work backwards. You have the density and you have the volume so plug them in. 19300kg/m^3=M/0.25 so multiply 19300 and 0.25 19300x0.25=? then uncube the answer to get final mass.

OpenStudy (aaronq):

For the first, you first have to find the volume using the formula for the volume a sphere, the 0.25 m is only the radius. \(V_{sphere}=\dfrac{4}{3}\pi r^3\) For the second, once you have the mass, you have to multiply the the acceleration due to gravity. \(\sf F=ma\)